### v2

# Rethinking the effects of data contamination in Code Intelligence

Abstract
In recent years, code intelligence has gained increasing importance in the field of automated software engineering. Meanwhile, the widespread adoption of Pretrained Language Models (PLMs) and Large Language Models (LLMs) has raised concerns regarding data contamination and its potential impact on model performance evaluation. This paper presents a systematic empirical study to investigate the fine-grained data contamination on code intelligence tasks. Our study involves diverse representative PLMs, namely RoBERTa and GPT-2, and LLMs, namely LLaMA and StarCoder, covering three major tasks: code translation, code generation, and code summarization. We categorize contamination scenarios into four types according to the code intelligence practice, namely input-only, output-only, unpaired, and paired contamination settings, and construct corresponding experimental and control groups for exploration.
Experimental results show that, under the pre-training, fine-tuning, and inference paradigm adopted by PLMs, even deliberately injecting paired contamination does not lead to significant performance overestimation. But direct inference or small-scale fine-tuning uncovers the contamination effects. In contrast, LLMs with pre-training and inference paradigm are significantly affected by the paired contamination. Apart from the above, other contamination scenarios have no impact on both PLMs and LLMs. Our findings challenge the conventional belief that contamination inevitably leads to performance overestimation, providing new insights into the evaluation and deployment of code intelligence models.
Index Terms:
data contamination, code intelligence tasks, large language model.I Introduction
In recent years, code intelligence has been a critical research area in the software engineering field, significantly improving software development and maintenance automation [1, 2, 3, 4, 5, 6]. This is mainly attributed to the advancement of Pre-trained Language Models (PLMs) [7, 8, 9, 10, 11, 12, 13, 14] and Large Language Models (LLMs) [15, 16, 17, 18, 19, 20] on code, especially their effective learning on massive Natural Language (NL) and Programming Language (PL) corpora. Despite their success, researchers began to question whether PLMs or LLMs’ high performances stem from their memorization of pre-training data, i.e., data contamination, and many related studies were proposed [21, 22, 23, 24, 25]. Nevertheless, most of them concentrate on sample-level contamination, which means both input and output portions are leaked in pairs in PLMs/LLMs’ pre-training corpora.
However, contamination does not always show up in samples, especially in the code intelligence field. For example, when doing code translation from Java to ArkTs, a declarative PL for application development in the HarmonyOS ecosystem [26]. PLMs/LLMs may not be pre-trained on ArkTS, but have seen massive Java code, including the one to be translated. In this case, a contamination scenario concerning only input portions occurs. Other scenarios are also prevalent in code tasks, such as the unit test case generation task. Perhaps both the focal methods and test cases have been pre-trained by PLMs/LLMs, but they are placed in different files on GitHub, leading to their unpaired learning during the pre-training stage of PLMs/LLMs. In this case, how does it affect PLMs/LLMs’ performance on downstream tasks? As such, this work seeks to fill this important gap.
Empirical Study: To explore the contamination effects in a fine-grained manner, we conduct a comprehensive empirical study involving two PLMs and two LLMs on three representative code intelligence tasks: code translation, code generation, and code summarization. In our analysis, we identify four typical types of data contamination scenarios commonly encountered in code intelligence: (1) input-only contamination, where only the input portion of the test sets appears in the pre-training corpora; (2) output-only contamination, where only the output portion is contaminated; (3) unpaired contamination, where both input and output components are seen in pre-training but not as matching pairs; and (4) paired contamination, where the exact input-output pairs from the test sets are found in the pre-training data of PLMs or LLMs.
As for experiments with PLMs, we adopt a pre-training, fine-tuning, and inference paradigm to simulate their practical usage. To circumvent uncontrolled data leakage, we perform pre-training from scratch [24, 27] and simulate contamination by injecting test data into the pre-training corpus according to different contamination settings. After fine-tuning on downstream tasks, we evaluate model performance on test sets and compare with the uncontaminated counterparts, thereby measuring the contamination effects. As for LLMs, we still follow their practical usage paradigm, i.e., pre-training then inference, to conduct the experiments. We use their publicly available pre-training corpora [16, 18] to extract appropriate samples according to different contamination settings as contaminated testing sets. As for uncontaminated counterparts, we design a series of perturbation rules to modify the contaminated testing sets, making them unseen to LLMs [28] while keeping sample complexities the same. All experiments are repeated five times for statistical analysis, thereby making comparisons and discussing the effects of contamination.
Main Findings: (1) PLMs with pre-training, then fine-tuning and inference paradigm are surprisingly not significantly affected by any contamination settings, overall yielding on average 0.16% and -0.06% of fluctuations on BLEU and METEOR over uncontaminated counterparts (p-value0.05). (2) Direct inference or small-scale fine-tuning uncovers the contamination effects on decoder-only PLMs, such as GPT-2, achieving 73.03% and 47.18% of overall lifting in terms of BLEU and METEOR (p-value0.05). (3) Large-scale fine-tuning tends to eliminate the contamination effects of decoder-only PLMs by overwhelmingly learning with underlying knowledge and task alignment specifications. (4) LLMs with pre-training and inference paradigm are only affected by paired contamination, obtaining consistent lifting of 10.87% and 5.67% on both metrics overall (p-value0.05). This paper makes the following contributions.
(1) To the best of our knowledge, this is the first systematic study exploring fine-grained data contamination effects in code intelligence.
(2) We propose a series of reproducible methods to extract contaminated samples for three code tasks and four contamination settings from the pretraining corpora of open-source LLMs, facilitating future contamination-related research.
(3) We conduct extensive experiments among diverse contamination settings, various PLMs and LLMs on multiple code intelligence tasks. A series of findings are revealed in this work, shedding light on future research and practice.
II Experimental Setup
We conduct the data contamination exploration on two mainstream models in code intelligence, i.e., Pre-trained Language Models (PLMs) and Large Language Models (LLMs). Towards each category, we experiment with multiple code-related tasks and varying contamination settings. In the following subsections, we first introduce our research questions, then present the corresponding explored models, tasks, and experimental methodologies.
II-A Research Questions
This paper proposes the following four research questions according to different contamination settings.
II-A1 RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
When constructing testing sets, only the input portion is included in the pre-training corpora of PLMs/LLMs, thereby simulating the input-only contamination. This is a prevalent scenario, the Java-to-ArkTS translation is a typical example as mentioned in Section I.
A prior study conducted by Jiang et al. [27] has claimed that the input-only and output-only contamination leads to performance overestimation for PLMs in Natural Language (NL) tasks, such as SQuAD and MMLU. However, NL tasks differ significantly from code intelligence tasks, and their experimental designs are flawed. First, the NL tasks carried out in Jiang et al.’s study only involve English text, leading to a relatively high similarity between their inputs and outputs compared with code intelligence tasks that always contain both NLs and code from different PLs. Second, the experiments of Jiang et al. only make comparisons in terms of average performance among only three trials without conducting statistically significant tests, making it hard to eliminate the influence of randomness and determine whether the outperformance of data contamination is consistent and prominent. Third, Jiang et al. only conducted their experiments on one PLM, i.e., GPT-2, without considering different architectures of PLMs and other LLMs. Therefore, the input-only contamination is highly worthwhile to re-evaluate in the code intelligence domain.
II-A2 RQ2: Does Output-Only Contamination Lead to Performance Overestimation?
Contrary to the above setting, output-only contamination places the output portion of testing sets in the pre-training corpora of PLMs/LLMs. Similar to the previous setting, this scenario is also practical in code intelligence and differs significantly from that in NL tasks. Thus, resolving this RQ is also necessary.
II-A3 RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
This is a kind of complete contamination of testing sets, but the input and output portions are not paired in the pre-training corpora of PLMs/LLMs. Recall the example we mentioned in Section I, when doing unit test case generation task, both focal methods and test cases may exist in the pre-training corpora of PLMs/LLMs, but their source files are split in the wild. Thus, these code samples will be pre-trained in PLMs/LLMs in a split manner. Considering the prevalence of this setting, unpaired input-output contamination is also meaningful and critical for exploration.
II-A4 RQ4: Does Paired Contamination Lead to Performance Overestimation?
As another kind of complete contamination, paired input-output contamination represents that both the input and output portions of testing sets have been seen by PLMs/LLMs, and they are concatenated in pairs during the pre-training stage. Many code intelligence tasks are very likely to involve this contamination, such as code generation, as Natural Language Descriptions (NLD) are always closely attached in front of corresponding code snippets, leading to their paired appearance in PLMs/LLMs’ pre-training. Theoretically, paired contamination is the most likely type to cause performance overestimation, and the LLM part has been investigated in [23, 22], but PLMs with pre-training, fine-tuning, and inference paradigm have not been explored.
II-B Models
II-B1 Pre-trained Language Models (PLMs):
Among this category, an encoder-only model, namely RoBERTa [29], and a decoder-only model, namely GPT-2 [30], are selected for experiments, as they have covered the most representative architectures of PLMs. We select RoBERTa with 125M parameters and GPT-2 with 150M parameters for experiments according to the load of our computational resources. To facilitate our research, we adopt one typical pre-training task for each PLM during its pre-training stage. RoBERTa is pre-trained with the Masked Language Modeling (MLM) [31] task, while the Causal Language Modeling (CLM) [32] task is used for GPT-2. Contamination injections are performed in the pre-training stage. Then, PLMs go through fine-tuning and inference stages to observe whether contamination induces performance overestimation compared with the uncontaminated counterparts.
II-B2 Large Language Models (LLMs):
As for this category, we select one general LLM, namely LLaMA [16], and one code LLM, namely StarCoder [18]. LLaMA is a family of multilingual LLMs with parameter sizes ranging from 7B to 65B. It is pre-trained on 1.4 trillion tokens of various open-source datasets, where 4.5% of the training base is code extracted from GitHub [33]. We select its 33B version for experiments. StarCoder is a Code LLM containing 15.5B parameters, trained on over 1 trillion code tokens. LLMs are experimented with in a pre-training then inference paradigm. Due to the massive scale and substantial training base of LLMs, retraining from scratch is impractical. We curate appropriate samples from their pre-training corpora for the contaminated testing set construction and experiments. The reasons we consider the above two LLMs are two-fold: First, they cover both general and code LLMs, making our exploration comprehensive. Second, they are two of the few open-source LLMs that also release their pre-training corpora, enabling us to construct contaminated testing sets and the uncontaminated counterparts from them.
II-C Code Tasks and Metrics
II-C1 Code Translation:
This task involves converting code from one PL to another while keeping the functionality the same, which is critical in scenarios such as migrating legacy systems and adopting modern frameworks. In this work, we focus on code translation between Java and C# for experiments. For simplicity, we use JavaC# to denote Java-to-C# translation while using C#Java to denote the opposite.
II-C2 Code Generation:
It is a process of automatically producing source code from a given NL requirement, aiming to reduce manual coding effort while ensuring correctness and efficiency. In this work, we adopt Java code generation for experiments, which is denoted as NLJava for simplicity.
II-C3 Code Summarization:
This task focuses on generating a NL description for a given code snippet, which is crucial for improving code readability, maintenance, and collaboration among developers. Still, we carry out experiments with code summarization in Java programs, which is denoted as JavaNL for simplicity.
II-C4 Metrics:
To automatically assess the quality of the outputs for the above code tasks, we adopt two widely used evaluation metrics, i.e., BLEU [34] and METEOR [35]. BLEU measures the precision of -gram overlap between the generated output and the reference, focusing on how much of the generated content is correct. In contrast, METEOR incorporates stemming, synonym matching, and alignment-based evaluation, placing more emphasis on recall—i.e., how comprehensively the generated output covers the information in the reference. By combining these two complementary metrics, we provide a more balanced and holistic assessment of generation quality from both precision and recall perspectives.
II-D Datasets
The following paragraphs present the datasets used in this paper for contamination construction, which has been partitioned into two parts (i.e., PLM and LLMs) for illustration.
II-D1 Dataset preparation for PLMs:
As for the pre-training corpus, we use the Java portion of the widely adopted CodeSearchNet dataset [36], which contains over 160,000 Java function snippets for mono-lingual pre-training. The downstream tasks contain code translation and code generation for the experiments of PLMs. To achieve this, we adopt the CodeTrans dataset [10], which includes 10,300 training samples, 500 validation samples, and 1,000 test samples for code translation between Java and C#. Besides, we use the NL-to-Java code generation dataset, namely CONCODE [37], for experiments, which contains 100,000 training samples and 2,000 validation and testing samples, respectively. Since we pre-train PLMs from scratch, we can use the testing sets of diverse code tasks to deliberately inject contaminated samples into pre-training corpora, also easily construct uncontaminated counterparts. Detailed experimental methods are elaborated in Section II-E.
II-D2 Dataset preparation for LLMs:
To construct contaminated datasets and their uncontaminated counterparts for LLMs, we curate their released pre-training corpora, i.e., Google BigQuery for LLaMA and The Stack (v1.2) for StarCoder. The GitHub dataset available on Google BigQuery constitutes 4.5% of the pre-training corpus of LLaMA, accounting for 328 GB. As the main training base of StarCoder, Stack (v1.2) consists of 6.4 TB of source code in 384 PLs. We construct our experimental data from these two corpora. In Section II-E, we elaborate on the detailed procedures to construct contaminated datasets and their uncontaminated counterparts according to the specific requirements of each contamination setting.
Model | Research Question | Language | Experimental Group | Control Group | Perturbance | ||||
Avg | Std | Med | Avg | Std | Med | count (ratio) | |||
LLaMA | RQ1 | Java (JavaC#) | 62.47 | 27.09 | 55.50 | 62.84 | 26.88 | 57.00 | 9.73 (15.33%) |
C# (JavaC#) | 58.86 | 28.18 | 51.00 | 59.63 | 28.23 | 51.50 | 9.27 (15.49%) | ||
NL (JavaNL) | 30.99 | 7.56 | 30.00 | 30.36 | 7.31 | 30.00 | 15.54 (48.58%) | ||
RQ2 | NL (NLJava) | 90.49 | 22.59 | 86.00 | 99.22 | 29.07 | 95.00 | 27.58 (30.15%) | |
RQ3 | Java (JavaC#) | 60.09 | 61.00 | 38.50 | 59.94 | 60.95 | 37.50 | 12.41 (20.31%) | |
C# (JavaC#) | 61.13 | 68.37 | 40.50 | 61.13 | 68.50 | 39.00 | 13.20 (21.25%) | ||
RQ4 | Java (NLJava) | 47.75 | 46.03 | 32.00 | 47.99 | 45.27 | 31.50 | 9.20 (18.87%) | |
NL (NLJava) | 63.70 | 46.04 | 47.5 | 63.94 | 46.47 | 48.00 | 12.32 (19.04%) | ||
Starcoder | RQ1 | Java (JavaC#) | 52.30 | 21.20 | 48.00 | 52.82 | 21.01 | 49.00 | 8.21 (15.70%) |
C# (JavaC#) | 52.01 | 22.53 | 49.00 | 52.58 | 22.44 | 49.00 | 7.86 (15.11%) | ||
NL (JavaNL) | 28.02 | 6.18 | 27.50 | 28.26 | 6.23 | 28.00 | 14.26 (15.70%) | ||
RQ2 | NL (NLJava) | 86.77 | 23.10 | 81.00 | 96.66 | 25.78 | 90.50 | 26.66 (30.72%) | |
RQ3 | Java (JavaC#) | 38.62 | 28.29 | 27.50 | 38.97 | 28.57 | 27.00 | 5.48 (14.19%) | |
C# (JavaC#) | 38.84 | 28.86 | 28.00 | 39.15 | 28.96 | 28.50 | 5.82 (14.98%) | ||
RQ4 | Java (NLJava) | 28.71 | 13.35 | 26.00 | 29.51 | 13.52 | 26.50 | 6.33 (22.05%) | |
NL (NLJava) | 45.02 | 22.99 | 38.00 | 45.90 | 23.28 | 39.00 | 10.46 (23.23%) |
Note: The Avg denotes the average number of tokens per sample, Std refers to the standard deviation of token counts, Med indicates the median number of tokens per sample, Language denotes the language of the input/output belongs to, and Perturbance count (ratio) represents the average number of perturbed tokens and their average proportion.
II-E Research Methodology
This section introduces the specific methodology that investigates the RQs proposed in Section II-A, especially the dataset construction for contaminated experimental groups and uncontaminated control groups.
II-E1 RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
In this RQ, PLMs are experimented on code translation and generation tasks, while LLMs are evaluated on code translation and summarization tasks. Accordingly, we construct a series of exclusive datasets for different code tasks and PLMs/LLMs, dividing them into contaminated experimental groups and uncontaminated control groups.
Pre-trained Language Models: For the code translation task (JavaC#) of PLMs, we dictate each Java code as input while using each corresponding C# code as output. As for the uncontaminated control group, we directly use our prepared 160,000 Java samples for PLMs to pre-train from scratch, then fine-tune and infer on the CodeTrans dataset. Towards the contaminated experimental group, for each sample in CodeTrans’s testing set, we inject its Java portion into the PLMs’ pre-training corpus while leaving its C# portion intact. To eliminate the bias of pre-training size brought about by the contamination injection, an equal number of randomly selected samples is removed from the pre-training corpus. As for the code generation task (NLJava) of PLMs, the inputs are NLDs, while the outputs are Java code snippets. We follow the same procedure for the code translation task to configure the experimental setting.
Large Language Models: For the code translation task (JavaC#) of LLMs, we first manually select 100 Java functions from both Google BigQuery’s GitHub and The Stack datasets as input portions of the contaminated experimental group. Afterwards, we manually migrate the above 100 Java functions into C#, forming code translation pairs, thereby ensuring the output portions are unseen to LLMs. To guarantee the migration quality, we follow Xue et al. [38]’s procedure and adapt their methodology into our manual migration, such as line-wise translation and keeping the naming conventions of C#. As for the code summarization task (JavaNL), we still adopt the above extracted Java functions as input portions, and manually craft their corresponding NLDs, thereby making the output portions (i.e., NLD) unseen to LLMs and forming the code summarization pairs of the contaminated experimental group. Similarly, we adhere to a series of standards when crafting NLDs. 1) NLDs must reflect the primary purpose of code snippets. 2) Limit NLDs to single sentences. 3) Focus on the high-level behaviors rather than low-level implementation details. Two authors with more than three years of Java and C# development experience are involved in the above manual procedures, one for implementation and another for double-checking.
Finally, to establish uncontaminated counterparts (i.e., control groups) in both code translation and summarization tasks, we perturb the above contaminated samples manually to ensure that both the input and output portions are unseen to LLMs. Specifically, the two authors participate in this procedure again. One is responsible for consistently perturbing both inputs and outputs, making sure they are still paired after perturbation. Besides, to guarantee that the task complexities are unchanged after perturbation for fair comparisons, only sub-token-level replacement is permitted (e.g., perturbing fieldName to fieldIdentifier), thereby maintaining that the code cyclomatic complexity [39] and code length are almost the same. After perturbation, another author double-checks the perturbation results to ensure the perturbation satisfies the consistency requirement mentioned above. To make readers learn more about our experimental datasets, we present Table I to illustrate their statistics, and a detailed illustration of each column is shown in the table note.
II-E2 RQ2: Does Output-Only Contamination Lead to Performance Overestimation?
In this RQ, both PLMs and LLMs are tested on code translation and generation tasks. The detailed experimental environment construction procedures are illustrated below.
Pre-trained Language Models: As for the code translation task (JavaC#) of PLMs, we follow all procedures in RQ1 to construct related experimental datasets. The only difference is that we use C# portion of CodeTrans’s testing set as contaminated data to inject into the pre-training corpora of both RoBERTa and GPT-2. As for the code generation task (NLJava), we still follow RQ1’s procedures except that we contaminate the Java portion (i.e., outputs) of CONCODE’s testing set.
Large Language Models: As for the experiments with LLMs, we take a different approach. Recall that we have prepared a series of input-output pairs for code translation and summarization tasks in RQ1. For the contaminated experimental group, their Java portions are extracted from Google BigQuery and Stack, which means they are contaminated. For simplicity, we directly swap the inputs and outputs for the contaminated experimental group and the uncontaminated control group, respectively, thereby constructing the output-only contamination setting. Thus, in Table I, statistics of C#Java of RQ2 are omitted. The above operations for the code translation task (C#Java) are enough owing to their line-wise equivalence. However, as for the code generation task (NLJava) derived from code summarization, to generate desired code snippets, previously crafted NLDs are too concise and lack implementation details. As such, we follow the criteria below to reconstruct NLDs for code generation. 1) Clearly state the input and output requirements. 2) Explicitly mention all external dependencies, such as libraries, classes, and interfaces. 3) Provide implementation-level details. Similarly, the previous two authors participate in this procedure again, one for implementation, another for double-checking. Specific statistics of experimental datasets are listed in Table I. As can be seen, the average token number of NL in RQ2 is larger than that in RQ1, owing to much more informative descriptions.
As for the uncontaminated control group, we follow the perturbation criteria mentioned in RQ1 to construct corresponding datasets, where their statistics concerning perturbation count and ratio are presented in Table I for reference.
II-E3 RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
In this RQ, we still focus on code translation and generation tasks for PLM. Towards LLM, two dual code translation tasks are carried out.
Pre-trained Language Models: To conduct experiments with PLMs in this setting, we followed all the procedures from RQ1 to construct the datasets. The difference lies in the fact that we not only include the input portion of downstream tasks, but also add the corresponding output part to the pre-training dataset. Specifically, for the code translation task (JavaC#), we inject both the Java and C# portions of CodeTrans’s testing set into the pretraining corpus of PLMs separately, and remove an equal number of randomly selected samples from the pre-training corpus to eliminate the impact of the difference in training amount. Similarly, this process is also applied to the code generation task (NLJava).
Large Language Models: For conducting the unpaired contamination experiment on LLMs, we need to build an entirely new testing set with both input and output portions explicitly existing in the pre-training corpora of LLMs as independent samples. Focusing on the code translation task, we can further swap their input and output portions to form two dual code translation tasks (JavaC# and C#Java). Specifically, we use tree-sitter to filter Java and C# functions from Google BigQuery and Stack, respectively. Then, considering programs of different PLs with the same method names normally also carry similar or even identical functionalities, we transform their method names to lowercase for exact matching, thereby obtaining over one hundred million pairs of programs with identical names. To reduce the manual workload, we calculate the BLEU score between the Java and C# code snippets in the collected pairs and remove those with BLEU scores lower than 0.6, as we hypothesize that the program pairs that could serve as mutual translations should have relatively high similarities. Subsequently, we filter out program pairs exceeding 15 lines to reduce the generation difficulty for the LLMs, alleviating experimental errors caused by overly long tasks. After performing these two steps of filtering, we conduct manual screening on the remaining program pairs with the previous two authors again, one for program pair selection and another for double-checking. After 48 hours of manual screening, we obtain 100 unpaired contaminated samples from Google BigQuery and Stack, respectively, thereby constructing the contaminated experimental group for LLMs in this setting. For the uncontaminated control group, we followed the same perturbation process as in RQ1 for construction. Finally, via swapping the inputs and outputs for experimental and control groups, respectively, we form two dual code translation tasks (JavaC# and C#Java) satisfying unpaired contamination settings. Detailed statistics are shown in Table I.
As for code generation and summarization tasks, it is challenging to find an independent NLD that coincidentally describes a code functionality but is not attached to a code snippet, because code and comments are normally closely attached in the wild. Therefore, we ignore the above two tasks for LLMs in the unpaired contamination setting.
II-E4 RQ4: Does Paired Contamination Lead to Performance Overestimation?
In this RQ, PLMs are still experimented with code translation and generation tasks. As for LLMs, we carry out the code generation task for exploration.
Pre-trained Language Models: To conduct code translation (JavaC#) experiments with PLMs in this setting, we still follow the procedures of RQ1 to establish the uncontaminated control group. As for the contaminated experimental group, we concatenate the input and output portions of CodeTrans’s testing set with a separator token <SEP> and inject these concatenated samples into the pre-training corpora of PLMs. We use <SEP> to separate different PLs, as it is a frequently-used method for bi-modal pre-training in practice [7, 12, 40]. Similarly, we also remove the corresponding number of samples from each pre-training corpus to eliminate the influence of pre-training amount differences. As for the code generation task (NLJava), we follow the same procedures for experiments.
Large Language Models: To construct paired contamination experiments on the code generation task (NLJava), we first use tree-sitter to identify and extract Java functions from both pre-training corpora. To curate samples containing both NLDs and code in order, we examine each sample to determine whether its NL description appears closely in front of the function signature. Afterwards, we perform manual filtering to select samples where the NL descriptions are accurate and informative enough to be used for code generation. The uncontaminated counterpart is prepared following the same perturbation procedures as RQ1.
As for code summarization, the inputs are code snippets while the outputs are NLDs. Although code-comment pairs are prevalent in practice, their comments are always in front of the corresponding code, which is opposite to the task definition of code summarization. Thus, we find difficulty curating such paired samples from LLMs’ pre-training corpora and ignore this task here. Moreover, in practice, programs of different PLs are very uncommon to be written in the same source files, even if they carry identical logic. Thus, it is hard to find two programs with equivalent logic pre-trained in pairs by LLMs. Therefore, we also ignore the code translation task for experiments with LLMs.
II-F Implementations
All experiments for PLMs and LLMs are conducted on two Nvidia A40 GPUs with 48GB of memory for each. For pre-training of the PLMs, we perform 20 epochs with a fixed learning rate of and a batch size of 32. For fine-tuning on downstream tasks, we train for 50 epochs on both the code translation and the generation tasks, ensuring full convergence. Our fine-tuning framework is adapted from the replication package of CodeXGLUE’s [10]. When inference, we fetch the first output for evaluation.
To enable direct inference by LLM, we deploy LLaMA and StarCoder locally with the model initialization and weight-loading from Hugging Face [41]. Subsequently, we perform inference on code tasks using a series of basic prompts.
To alleviate the chaos of outputs, we additionally provide a one-shot example for in-context learning and retrieve only the first output for evaluation. The inference process is conducted with the following decoding parameters: temperature = 0.8, top_p = 0.95, and top_k = 50.
To mitigate the influence of randomness and enhance the reliability of the results, all experiments concerning PLMs and LLMs are repeated five times, respectively. Besides, we employ the non-parametric Mann–Whitney U test [42] with a two-sided hypothesis to assess whether the performance differences between experimental and control groups are statistically significant. For each test, the null hypothesis is defined as: “There is no statistically significant difference between the results of the two groups.” All statistical tests are conducted at a significance level of = 0.05, corresponding to a 95% confidence level for determining statistical significance.
III Experimental Result
This section discusses the experimental results targeting each RQ proposed in Section II-A. Owing to the page limit, we only showcase results with their average score and standard deviation.
III-A Results of the Input-Only Contamination Setting (RQ1)
As illustrated in Section II-A1, we conduct experiments on code translation and generation tasks for PLMs while carrying out code translation and summarization tasks for LLMs. Below, we introduce their experimental results and corresponding analysis, respectively.
As for PLMs, the performance of RoBERTa and GPT-2 under control and experimental groups is reported in Table II. Apparently, the performance in the contaminated experimental group does not exhibit a significant performance gain compared to the uncontaminated counterparts, regardless of the code translation or generation task and RoBERTa or GPT-2. For example, on the code translation task with GPT-2, the experimental group’s average BLEU score is only 0.97% higher than that of the control group, which falls within the margin of experimental error. Even surprisingly, the model performance of the control group surpasses that of the experimental group when experimenting with RoBERTa, although with only 0.046%–0.86% across both metrics and code tasks. Statistical tests further support this observation, with -values exceeding 0.05 among five consecutive trials, indicating no significant differences. These results suggest that the performance of PLMs will not be overestimated in the downstream tasks, even though their pre-training corpora are contaminated by the input portions of testing sets. This conclusion aligns with our expectation and is reasonable for explanation, as PLMs do not learn any mapping from input to output portions of the testing sets. Their pre-training corpora only contain questions but no answers.
As for LLMs, their experimental results of the input-only contamination setting are also listed in Table II. As can be seen, the inference performance in the contaminated experimental group and the uncontaminated control group has no obvious discrepancy among the five consecutive trials. Besides, statistical tests further verify the nonsignificance of this difference with p-values all larger than 0.05. The reason is the same as that we mentioned in the PLM part, as LLMs only saw the input portions during their pre-training, considering the substantial discrepancies between inputs and outputs in code tasks, by no means can they carry any superiority in generating the output portions.
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.296 | / / 0.531 |
NLJava | / / 0.210 | / / 0.095 | |
GPT-2 | JavaC# | / / 0.296 | / / 0.249 |
NLJava | / / 0.210 | / / 0.296 | |
LLaMA | JavaC# | / / 0.676 | / / 0.296 |
JavaNL | / / 0.531 | / / 0.210 | |
Starcoder | JavaC# | / / 0.296 | / / 0.403 |
JavaNL | / / 0.531 | / / 1.000 |
Note: In the Task column, the contaminated part is highlighted in bold. The results are presented in the form of , where is the average result of 5 experiments, and is the corresponding standard deviation. p-value specifies the significance of statistical tests. We highlight the higher scores between the two groups with a yellow background. Subsequent tables follow the same convention.
III-B Results of the Output-only Contamination Setting (RQ2)
As elaborated in Section II-A2, we carry out experiments on code translation and generation tasks for both PLMs and LLMs. Below, we introduce their experimental results and corresponding analysis, respectively.
As for PLMs, the results for RoBERTa and GPT-2 under this setting are shown in Table III. Similar to the input-only contamination experiments, there is no significant difference in performance between the contaminated experimental group and the uncontaminated control group, regardless of any code intelligence task and PLMs under test. Specifically, experiments with RoBERTa all demonstrate a slightly superior performance in the contaminated scenarios, while GPT-2 performs the opposite. However, the differences are minimal. For example, in terms of BLEU score on the code translation task, the experimental group of Roberta was only 0.21% higher than the control group, and the control group of GPT-2 was just 0.51% higher than the experimental group on average. Besides, when focusing on statistical tests, their results among the five trials do not manifest any statistical significance with p-values all larger than 0.05. This indicates that exposure to the output portions of the testing set during the pre-training stage does not cause performance overestimation for PLMs.
As for LLMs, the inference results under the output-only contamination setting are presented in Table III as well. The observed performance distribution aligns with that of PLMs, i.e., neither the contaminated experimental group nor the uncontaminated control group dominate on code intelligence tasks, and none of the statistical tests demonstrate significant differences with p-values larger than 0.05 consistently. This suggests that contamination appearing on output portions of testing sets will not induce performance overestimation in downstream tasks.
In summary, neither PLMs nor LLMs will be influenced by the output-only contamination in their pre-training corpora. As we explained in Section III-A, models only saw the partial samples of testing sets, which is not helpful for them to infer the rest of the samples, as the differences between input and output portions of code intelligence tasks are substantial.
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.404 | / / 0.095 |
NLJava | / / 0.144 | / / 0.676 | |
GPT-2 | JavaC# | / / 0.531 | / / 0.676 |
NLJava | / / 0.600 | / / 0.403 | |
LLaMA | C#Java | / / 0.676 | / / 1.000 |
NLJava | / / 0.403 | / / 0.210 | |
Starcoder | C#Java | / / 0.144 | / / 0.403 |
NLJava | / / 0.210 | / / 0.095 |
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.210 | / / 0.296 |
NLJava | / / 0.531 | / / 0.144 | |
GPT-2 | JavaC# | / / 1.000 | / / 0.676 |
NLJava | / / 0.835 | / / 0.531 | |
LLaMA | JavaC# | / / 0.144 | / / 0.095 |
C#Java | / / | / / 0.060 | |
Starcoder | JavaC# | / / 0.676 | / / 1.000 |
C#Java | / / | / / 0.095 |
Note: Hypothesis test values less than 0.05 are indicated in red. Subsequent tables follow the same convention.
III-C Results of the Unpaired Contamination Setting (RQ3)
As presented in Section II-A3, unpaired contamination experiments are conducted on code translation and generation tasks for PLMs. As for LLMs, we carry out the code translation task solely for exploration. Below, we introduce and analyse the experimental results of PLMs and LLMs, respectively.
Similarly, we present the PLMs’ results first, which are shown in Table IV. It should be noted that both the inputs and outputs are highlighted in bold under the Task column, as both of them are contaminated in the experimental group. As can be seen, although PLMs in the contaminated experimental groups perform slightly better than those in the uncontaminated control groups, their superiority is very limited, ranging from 0.06%–0.52% in terms of BLEU while 0.11%–0.68% in terms of METEOR, and the statistical tests still demonstrate that there is no significant difference between the two groups with p-value larger than 0.05. This suggests that merely observing both inputs and outputs independently is insufficient for the PLMs to learn their correspondence in downstream tasks, as the input-output mappings are not constructed during the pre-training stage.
The experimental results of the unpaired contamination setting for LLMs are shown in Table IV, where the code translation task is conducted in both Java-to-C# and C#-to-Java translation. As can be seen, LLMs perform slightly better in the contaminated experimental group than in the uncontaminated control group across most of the translation scenarios and evaluation metrics. Nonetheless, their performance differences still cannot be considered significant, as the p-values of the statistical tests are larger than 0.05 in most cases, except the C#-to-Java translation when evaluating in the BLEU score, which has been highlighted in red. Nevertheless, their performance in terms of METEOR does not demonstrate statistically significant differences between the two groups. Besides, theoretically, unpaired contamination only involves pre-training with both input and output portions of testing sets, but the mapping between them is unlearned. Therefore, we can only infer that this is an experimental coincidence rather than contamination-induced performance overestimation.
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.531 | / / 0.676 |
NLJava | / / 0.835 | / / 0.676 | |
GPT-2 | JavaC# | / / 0.296 | / / 0.144 |
NLJava | / / 0.403 | / / 0.676 | |
LLaMA | NLJava | / / | / / |
Starcoder | NLJava | / / | / / |
III-D Results of the Paired Contamination Setting (RQ4)
As illustrated in Section II-A4, paired contamination experiments are carried out on code translation and generation tasks for PLMs, while for LLMs, only the code generation task is experimented with. Below, we introduce and analyse PLMs and LLMs’ experimental results, respectively.
The experimental results of PLMs are shown in Table V. As can be seen, neither the contaminated experimental group nor the uncontaminated control group manifests consistent superiority in either code translation or generation tasks and either RoBERTa or GPT-2. Besides, the statistical tests also demonstrate non-significant differences with all p-values larger than 0.05 among five consecutive trials. The performance of RoBERTa is reasonable, as its pre-training task is MLM, which is dictated to learn to infill the masked tokens during the pre-training stage and is not aligned with the inference procedure of any downstream tasks. Besides, to implement the generative task, RoBERTa need to be appended with a decoder structure, further weakening the impact of pre-training contamination on downstream tasks. On the contrary, the experimental results of GPT-2 are kind of surprising and opposite to our expectations, because it is an auto-regressive model [32] and was pre-trained by the CLM task from scratch in our experiments. Considering that the input and output portions of the testing sets are contaminated in the pre-trained corpus of GPT-2, it should have learned how to generate correct tokens during the inference stage, theoretically. But the fact is not the case. We further explore possible explanations in Section IV.
As for LLMs, their experimental results are shown in Table V as well. Apparently, both LLMs’ performance in contaminated experimental groups significantly surpasses that in uncontaminated control groups. Specifically, the outperformance reaches 14.34% in terms of BLEU and 7.26% in terms of METEOR for LLaMA. As for StarCoder, the overestimation can be 7.40% and 4.07% in terms of each metric in order. Moreover, statistical tests with p-values smaller than 0.05 further testify to the significance of the difference. Considering that LLMs under test are all decoder-only architectures and are pre-trained with either CLM (i.e., LLaMA) or Fill-in-the-Middle (FIM) task (i.e., StarCoder), the explanation is straightforward. Because a decoder-only architecture implies its auto-regressive nature in practical usage. Besides, their pre-training tasks help them learn to generate subsequent tokens given prefixes, and even the FIM task broke away from the limitation of left-to-right generation [43]. Both of the above reasons make LLMs’ pre-training stage almost or exactly match their inference stage. As such, paired contamination leads to performance overestimation of LLMs in downstream code intelligence tasks.
IV Discussion
As stated in RQ4, although we inject paired contaminated data into the pre-training stage of GPT-2, the overall model performance did not exhibit a significant improvement, even though it is an auto-regressive model and its pre-training task is almost the same as LLMs under test. The conclusion is completely opposite. Considering there is a fine-tuning stage between pre-training and inference for PLMs, we propose our RQ5 naturally: Whether the expected contamination effects appear if we drop the fine-tuning stage from experiments of GPT-2, and how fine-tuning influences the contamination effects?
To do this, we conduct a direct inference experiment using GPT-2 with previously pre-trained models in RQ4. The experimental results are shown in Table VI. We find that the model performance in the contaminated experimental group significantly outperforms that of the uncontaminated counterparts, where the former achieves 26.52% and 20.57% of superiority in terms of BLEU and METEOR on Java-to-C# translation, respectively. As for the code generation task, contamination even brings about overestimation by 119.53% and 73.79% in terms of each metric in order. Besides, statistical tests also demonstrate the significance of the performance differences with p-values all smaller than 0.05. This indicates that contamination indeed affects GPT-2, making its performance overestimated with the pre-train-then-inference paradigm, but this effect was covered up by the subsequent fine-tuning. As mentioned in Section II-D and II-F, our fine-tuning was conducted with 10,300 and 100,000 training samples for code translation and generation tasks, respectively. Both are trained for 50 epochs. Thus, we infer that fine-tuning brings about abundant underlying knowledge and task alignment specifications [44], making the effects of contamination in the pre-training stage negligible.
To further investigate the influence of fine-tuning on contamination effects, towards the code translation and generation tasks, we reduce the fine-tuning size to 500 samples and progressively enlarge the training epochs (E=0, 1, 5, 10, 25, 50) for multiple experiments, thereby controlling the learning of underlying knowledge and task alignment specifications. As with our previous RQs, we repeat the experiment 5 times for each setting, and the results are presented in Fig. 1, where epoch = 0 denotes zero-shot inference without fine-tuning. For each pair of control and experimental groups, we use a black line to connect them, thereby making their performance differences more clear. Meanwhile, p-values for each experiment are plotted with red dots connected by dashed lines in Figure 1. As can be seen, the contaminated experimental group performs significantly better than the uncontaminated counterpart initially, i.e., E10, for code translation while E5 for code generation. With the training epochs enlarging, which means more underlying knowledge and task alignment specifications are learned by GPT-2, the contamination effects gradually eliminate. Besides, we also find an interesting phenomenon: the performance of GPT-2 in both experimental and control groups shows a trend of declining first, then rising across both code tasks. A potential explanation is that the pre-training stage of our experiments mainly learn to generate Java code given Java prefixes, which is relatively different from the fine-tuning tasks, i.e., code translation and generation. Thus, there are certain conflicts between them, leading to the re-learning of the task alignment on both downstream code tasks at the initial fine-tuning stage. With the training continuing, GPT-2 gradually overcomes the task alignment conflicts, thereby recovering and even improving its performance.
Task | BLEU (Experimental / Control / p-value) | METEOR (Experimental / Control / p-value) |
JavaC# | / / | / / |
NLJava | / / | / / |
V Implications
This paper reveals the impact of various data contamination settings on the performance of PLMs and LLMs in practical usage. Based on the experimental results, we summarise the following implications.
V-A Implications for Researchers
This paper serves as the first systematic study delving into the practical effects of different contamination settings on PLMs and LLMs. To achieve this, we design rigorous experiments for both PLMs and LLMs. In particular, we propose a series of efficient and reproducible methods to extract contaminated samples from tremendous pre-training corpora of LLMs and construct uncontaminated counterparts applicable for diverse contamination settings, which can be reused or refined by future related research. Secondly, our research concluded that contamination effects do not appear in all scenarios and all models. Researchers can conduct their LLM/PLM-related experiments according to our conclusions to deal with the data contamination problem more precisely.
V-B Implications for Practitioners
Based on our extensive experimental results, we provide guidance for developers to better understand the impact of data contamination in code intelligence. Firstly, we find that although PLMs suffer from contamination during pre-training, their performance does not show significant improvement on the test set after fine-tuning, indicating that the impact of contamination can be effectively mitigated to some extent. Therefore, deploying PLMs with fine-tuning in practice is safe for data contamination, contrary to the conclusion from NLP research [45]. Therefore, contamination in fine-tuning should be given more attention instead of pre-training for PLMs. Secondly, practitioners must exercise heightened vigilance around paired contamination for the usage of LLMs, as this configuration makes model performance significantly overestimated. This differential impact between PLMs and LLMs necessitates tailored contamination mitigation strategies in practice.
VI Threats to Validity
Limited contamination scale: In this study, the proportion of contaminated data in experiments of both PLMs and LLMs is relatively small compared to their pre-training corpora, without further exploration of the larger scale of contamination. However, it is worth noting that in real-world scenarios, benchmark evaluation datasets, such as HumanEval [15] and ClassEval [46], typically constitute only a tiny fraction compared to the scale of the PLM/LLMs’ pre-training corpora, which aligns with our experimental settings. Therefore, our experimental results remain reasonably credible for code intelligence practice despite this potential limitation.
Perturbation quality: when constructing uncontaminated datasets for LLMs, we manually perturb the contaminated data to be unseen to LLMs, leading to two threats concerning data quality, i.e., 1) whether the perturbed sample can be considered uncontaminated, 2) whether the perturbation affects the sample complexities. To achieve this, we adopt a relatively high perturbation rate of 14.19%–48.58% on average across various experiments (as shown in Table I), thereby making uncontaminated samples as different from the original versions as possible, thereby alleviating the former threat. Besides, two authors of this paper participate in the perturbation with a double-checking mechanism to ensure the consistency of token lengths and cyclomatic complexity before and after perturbation. Therefore, the threat of the latter is also minimised.
VII Related Work
VII-A Code Intelligence Tasks on PLMs and LLMs
Recent years have witnessed extensive research on code intelligence tasks based on PLMs/LLMs, such as code generation [47, 48], code summarization [49, 50], and code translation [51, 38]. Early efforts, such as CodeBERT [52] and GraphCodeBERT [8], jointly modeled NL and PL, enhancing semantic understanding of code through contrastive learning and structural awareness. Subsequently, PLBART [11], based on an encoder-decoder architecture, demonstrated strong transfer capabilities. CodeT5 [12] further pre-trained on T5 [53], achieving state-of-the-art performance on a variety of code understanding and generation tasks. On the LLM front, Codex [15] and CodeGen [17] have shown remarkable capabilities in multi-turn reasoning and complex programming scenarios. More recent works, such as StarCoder [18] and WizardCoder [54] explore the upper performance bounds of open-source LLMs in code intelligence, emphasizing the importance of model scale, instruction tuning, and multilingual training for improving code comprehension and generation abilities.
VII-B Studies On Data Contamination
Data contamination studies have been investigated for years, many are dedicated to contamination detection. For instance, Golchin et al. [25] proposed both instance-level and partition-level contamination detection, while Dong et al. [55] proposed CDD, using output distributions to detect contamination in LLMs on both code and NLP tasks. Deng et al. [56], instead, proposed TS-Guessing to detect contamination by examining whether LLMs can produce masked wrong answers.
As for research exploring contamination effects, early research was conducted in the NLP field. For example, Magar et al. [24] explored to what extent BERT exploits or just memorises the contaminated data for downstream tasks. Jiang et al. [45] explored the input-only/output-only contamination effects on GPT-2. Subsequently, contamination has also been studied in the code intelligence field, Riddell et al. [23] and Ramos et al. [22] quantified the contamination effects on LLM-based code generation and bug fixing, respectively. Cao et al. [21] extracted data before and after LLMs’ cut-off date to study the effect of countermasures on coding tasks.
Nonetheless, owing to the distinctive features of code tasks, as we mentioned in Section II-A1, different conclusions from the NLP research concerning input/output-only contamination are obtained. Focusing on the domain of code intelligence, most studies only consider sample-level contamination [21, 23, 22], i.e., the fourth setting in this paper. As such, the effects of other fine-grained contamination issues remain unanswered questions, which is a striking gap this paper fills.
VIII Conclusion
This paper presents a comprehensive empirical study to reveal the impact of different data contamination scenarios on code intelligence. The whole experiment covers diverse PLMs/LLMs, and code tasks, uncovering different conclusions that were released before in the NLP field, and challenging the conventional belief that contamination inevitably induces performance overestimation in the code intelligence field, manifesting the necessity of our research. Future work could further explore the effects of large-scale contamination and validate the generality of the findings across tasks and PLs.